OpenStudy (anonymous):

Solve the equation is the answer 9? \[\sqrt{x-7}-2=0\]

OpenStudy (anonymous):

should be 11

OpenStudy (anonymous):

\[\sqrt{x-7}=2 \rightarrow x-7=4 \rightarrow \]

OpenStudy (anonymous):

11

OpenStudy (anonymous):

how are you getting the 4 part? i see 4+7 = 11

OpenStudy (anonymous):

\[\sqrt{x-7}-2=0\]\[\sqrt{x-7}=2\]\[(\sqrt{x-7})^2=2^2\]\[x-7=4\]\[x=4+7\]\[x=11\]

OpenStudy (anonymous):

ok i think i am following, i hope

OpenStudy (anonymous):

You most likely forgot to square both side when getting rid of the sr root

OpenStudy (anonymous):

i did forget, i think that is where i had it screwed up

OpenStudy (anonymous):

yup